# Matbench Discovery A framework to evaluate machine learning crystal stability predictions

### v2

@@ 3 Discussion
We have demonstrated the effectiveness of ML-based triage in HT materials discovery and posit that the benefits of including ML in discovery workflows now clearly outweigh the costs. LABEL:{tab:metrics-table-uniq-protos} shows in a realistic benchmark scenario that several models achieve a discovery acceleration greater than 2.5 across the whole dataset and up to 5 when considering only the 10k most stable predictions from each model (table 3). When starting this project, we were unsure which is the most promising ML methodology for HT discovery. Our findings demonstrate a clear superiority in the accuracy and extrapolation performance of UIPs like CHGNet, M3GNet and MACE. Modeling forces enables these models to chart a path through atomic configuration space closer to the DFT-relaxed structure from where a more informed final energy prediction is possible. The consistently linear log-log learning curves observed in related literature [46] suggest that further decreases in the error of UIPs can be readily unlocked with increased training data.
Despite its fast pace of progress, the field of universal ML potentials is still in its early stages. To realize the full potential of these models, we will need to pivot significant resources to generate large quantities of higher-than-PBE fidelity training data. That said, we believe our work (e.g. fig. 8) and others [27] have caught early glimpses that sufficiently large training sets allow us to narrow the performance gap between in-domain and out-of-domain performance. This should further increase the discovery acceleration factor, resulting in a virtuous cycle between ML models accelerating the creation of new training data while becoming more accurate and even stronger accelerators after retraining. Meanwhile, the setup cost and learning curve of ML potentials are steadily decreasing, likely with a much lower floor than DFT given the intricacies of HT ab-initio simulation. Combined with their already strong performance, this suggests that UIPs combined with data accumulation over the next few years may unlock our way to a comprehensive yet cheap map of the PES that truly commoditizes HT materials discovery.
Despite this optimistic outlook, not only has the predictive power UIPs attained for crystal stability over the last 3 years since [13] sobering 2020 analysis yet to enter common knowledge, but many remaining shortcomings invite further research. For example, all models are biased towards stability, overestimating the correct number of stable materials in our test set anywhere between 30% and 133%. Moreover, even the best models in our analysis still exhibit a significant fraction of false positive predictions, even for materials as high as above the hull. This indicates the need to train future models on even more diverse datasets that are less biased towards ground state crystals than the MP training set we used. No such dataset built with MP-compatible DFT settings exists to our knowledge, inviting future efforts in this direction.
@@
Looking beyond mere thermodynamic stability prediction at zero Kelvin for the purpose of materials discovery, future materials science will also require understanding and predicting the properties under varying environmental conditions such as finite temperature and pressure. This is where interatomic potentials can unlock further utility. Their force predictions enable insights into dynamical properties. One open question of particular relevance is the extent to which such models may aid research into the computational prediction of synthesis pathways. Many existing approaches for reaction pathway prediction involve the use of heuristic rules for dealing with the significant added complexity of metastability alongside traditional ground state ab-initio data [47, 48, 49]. These algorithms stand to benefit massively from more efficient estimates of the reaction energy barriers that future UIPs may provide. Provided they reach sufficient accuracy, this may enable timely calculation of approximate reaction barriers [50] and open up a whole new field to high-throughput inquiry.
@@ 4 Acknowledgments
We thank Rickard Armiento, Felix A. Faber and Abhijith S. Parackal for helping develop the evaluation procedures for Wren upon which this work builds. We also thank Rokas Elijosius for assisting in the initial implementation of Wrenformer.
@@ 5 Author Contributions
Janosh Riebesell: Methodology, Software, Data Curation, Training and Testing Models, Formal Analysis. Rhys Goodall: Conceptualization, Software, Formal Analysis. Philipp Benner: Software, Training ALIGNN. Yuan Chiang: Training MACE. Bowen Deng: Curating MPtraj dataset and training CHGNet. Alpha Lee: Supervision. Anubhav Jain: Supervision. Kristin Persson: Funding Acquisition.
@@ 7 Data availability
We chose the latest Materials Project (MP) [3] database release (v2022.10.28 at time of writing) as the training set and the WBM dataset [29] available at https://figshare.com/articles/dataset/22715158 as the test set for this benchmark. A snapshot of every ionic step including energies, forces, stresses and magnetic moments in the MP database is available at https://figshare.com/articles/dataset/23713842.
@@ Appendix A Supplementary Information
A.1 Metrics for 10k materials predicted most stable
@@
Unlike table 1 which evaluates model performance on the subset of unique WBM structure prototypes, table 2 includes the full WBM test set of 257k materials. The 44k additional materials in table 2 comprise discarded earlier due to having a matching structure prototype in MP plus materials which are prototype duplicates of another WBM material with lower energy. The most noteworthy difference between the two tables is a drop in DAF for all models in table 2 compared to table 1. MACE for example achieves a DAF of 3.5 on the full test set (table 2) compared to 3.85 on the subset of 215.5k materials with unique prototypes (table 1). This 10% increase is largely due to a 10% decrease in the fraction of materials below the MP convex hull: 15.3% (32,942 out of 215,488) in table 1 vs 16.7% (42,825 out of 256,963) in the full dataset (table 2). Since DAF is the ratio of model’s precision when predicting stability to the prevalence of stable structures, a lower prevalence results in a higher DAF.
@@
While prototypes are non-trivial to match, thus potentially introducing bias into the test set by trying to deduplicate them, we still opted to feature table 1 in the main text since the removal of overlapping prototypes with MP makes it more closely reflect a model’s true ability to extrapolate to out-of-domain materials. Prototypes were matched based on an Aflow-style [51] Wyckoff representation encoding the crystal’s prototype as implemented in aviary.wren.utils.get_aflow_label_from_spglib| [10]. This string is invariant to relaxation (unless the system changes spacegroup during relaxation, usually by gaining symmetry) and will count structures with different lattice parameters as duplicates if both are expected to relax to the same ground state.
@@
A real-world discovery campaign is unlikely to validate all stable predictions from a given model as we did in LABEL:{tab:metrics-table-uniq-protos}. Presumably, it will rank model predictions from most to least stable and follow that list as far as time and compute budget permits. Assuming that increases in compute resources will allow average future discovery campaigns to grow in scope, we believe 10k model validations to be a reasonable scope for average campaigns. This is what table 3 simulates by calculating classification and regression metrics for the 10k test set materials predicted to be most stable by each model. We again show dummy performance in the bottom row. Note that each model is now evaluated on a different slice of the data. However, the bottom row still shows dummy performance across the whole dataset. CHGNet and M3GNet achieve an impressive precision of 0.83 and 0.8, respectively. In concrete terms, this means in a discovery campaign that validates 10 k model predictions from a search pool of 257 k crystals that are chemically dissimilar from the training set and of which 0.167 are stable, CHGNet and M3GNet would deliver 4 stable structures for every 5 predictions validated. We emphasize that in light of the significant resulting increase in stability hit rate, these models are well worth integrating into future materials searches.
A.2 ROC Curves
@@
A material is classified as stable if the predicted is below the stability threshold. Since all models predict (and M3GNet predicted energies are converted to formation energy before stability classification), they are insensitive to changes in the threshold . The receiver operating characteristic (ROC) curve for each model is plotted in fig. 3. M3GNet wins in the area under curve (AUC) with 0.87, coming in 1.34x higher than the worst model, the Voronoi Random Forest. The diagonal ‘No skill’ line shows the performance of a dummy model that randomly ranks material stability.
A.3 Parity Plots
@@
Figure 4 shows that all models do well for materials far below the convex hull (left side of the plot). Performance for materials far above the convex hull is more varied with occasional underpredictions of the energy of materials far above the convex hull (right side). All models suffer most in the mode of the distribution at .
@@
Wrenformer is the only model with a large number of severe energy overpredictions at along the positive axis. We investigate these failure cases in more detail in fig. 5 and find these overpredictions to be dominated by spacegroup 71 with poor representation in the training data.
@@
Beyond these MACE outliers visible in the plot, MACE exhibits another rare but reproducible type of failure case, in which the final predicted energy after relaxation is off by several orders of magnitude. The largest ‘derailed’ prediction was eV/atom for wbm-3-31970 (formula HIr). In each case, the MACE relaxation exhausted the maximum number of ionic steps set to 500 and caused a volume implosion from initial cell volumes of hundreds to relaxed cell volumes of tens of cubic Angstrom. Using the checkpoint trained on the M3GNet dataset which we received from the MACE authors, this failure mode occurred for several hundred of the 250k test set crystals. Using the checkpoint we trained ourselves on the MPTrj dataset, it affects only 44 test crystals, suggesting that these holes in the MACE PES can perhaps be fully plugged by further increasing the training set or even changing the loss function. Further analysis is ongoing. Since these derailed values are easily identified in practice when actually performing a prospective discovery campaign, we excluded them from the MACE parity plat and all other downstream analyses.
A.4 Spacegroup prevalence in Wrenformer failure cases
Figure 5 shows 456 + 194 (70%) of the failure cases in the shaded rectangle are two prototypes in spacegroup 71. The occurrence of those same prototypes in the MP training set shows almost no data support for the failing prototypes. This suggests the reason Wrenformer fails so spectacularly on these structures is that it cannot deal with structure prototypes it has not seen at least several hundred examples of in its training data. This suggests that there are stronger limitations on how much the discrete Wyckoff-based representation can extrapolate to new prototypes compared to the smooth local-environment-based inputs to GNN-type models.
A.5 Hull Distance Box plot
BOWSR has the largest median error, while Voronoi RF has the largest IQR. Note that MEGNet and CGCNN are the only models with a positive median. Their hull distance errors are biased towards more frequently predicting thermodynamic instability, explaining why they are closest to correctly predicting the overall number of stable structures in the test set (see fig. 1).
A.6 Classification Histograms using Model-Predicted Energies
Note the CGCNN+P histogram is more strongly peaked than CGCNN’s which agrees better with the actual DFT ground truth distribution of hull distances in our test set. This explains why CGCNN+P performs better as a regressor, but also reveals how it can perform simultaneously worse as a classifier. By moving predictions closer to the stability threshold at above the hull, even small errors are significant enough to tip a classification over the threshold.
A.7 Measuring extrapolation performance from WBM batch robustness
@@
Figure 8 shows the rolling MAE as a function of distance to the convex hull for each of the 5 WBM rounds of elemental substitution. These plots show a stronger performance decrease for Wrenformer and Voronoi RF than for UIPs like M3GNet, CHGNet, MACE and even force-less GNNs with larger errors like MEGNet and CGCNN.
MEGNet and CGCNN both incur a higher rolling MAE than Wrenformer across all 5 batches and across most or all of the hull distance range visible in these plots. However, similar to the UIPs, MEGNet and CGCNN exhibit very little degradation for higher batch counts. The fact that higher errors for later batches are specific to Wrenformer and Voronoi RF suggests that training only on composition and coarse-grained structural features (spacegroup and Wyckoff positions in the case of Wrenformer; coordination numbers, local environment properties, etc. in the case of Voronoi RF) alone is insufficient to learn an extrapolatable map of the PES.
@@
A.8 Largest Errors vs. DFT Hull Distance
@@
Given the large variety of models tested, we asked whether any additional insight into the errors can be gained from looking at how the predictions vary between different models. In fig. 9 we see two distinct groupings emerge when looking at the 200 structures with the largest errors. This clustering is particularly apparent when points are colored by model disagreement.
A.9 MEGNet formation energies from UIP-relaxed structures
Interestingly, table 4 shows that MEGNet RS2RE performs worse than CHGNet and M3GNet even when it has access to the ground truth DFT relaxed structures when making relaxed energy predictions. While the regression and classification metrics for MEGNet RS2RE are better than M3GNet→MEGNet and CHGNet→MEGNet, the difference is surprisingly small, especially compared to the regression performance of standalone MEGNet. This suggests that for downstream use, at least by an imprecise method such as another ML model, the UIP-relaxed structures are almost as helpful as the DFT-relaxed ones.
CHGNet, M3GNet and MEGNet were all trained on slightly different targets. While MEGNet was trained to predict formation energies, M3GNet and CHGNet were trained on raw DFT energies with the important distinction that CHGNet targets include MP2020 energy correction scheme [52] while M3GNet targets do not.
Changing the target from raw DFT energies to formation energies in principle merely constitutes a change of gauge since the difference is just a linear transformation of subtracting the elemental reference energies weighted by composition. But in practice, one might expect the problem of stability prediction to be easier when trained on formation energies, as the model output is one less step removed from the final target of convex hull distance. However, empirical evidence so far suggests the opposite. Both UIP→MEGNet combos perform worse than those same UIPs by themselves. There are too many confounding effects at play to draw firm conclusions but this tentatively suggests using formation energy as targets offers no benefits over raw DFT.
A.10 Exploratory Data Analysis
To give high-level insights into the MP training and WBM test set used in this work, we include element distributions for structures in both data sets (fig. 11). To show how frame selection from MP structure relaxation affected relative elemental abundance between MP relaxed structures and the snapshots in MPtrj, fig. 12 shows element occurrence ratios between MPtrj and MP. Believing MPtrj to be an influential dataset for the near-term continued development of universal interatomic potentials, we plot histograms showing the distributions of target values for energies, forces, stresses and magnetic moments in fig. 14. Similarly, fig. 15 plots the distribution of formation energies and convex hull distances (with respect to the convex hull spanned by MP materials only) of the WBM test set. This plot not only gives insight into the nature of the dataset but also emphasizes the increased difficulty of stability vs formation energy prediction arising from the much narrower distribution of convex hull distances compared to the more spread-out formation energy distribution.
To get a visual sense of coverage of atomic configuration space afforded by WBM compared to MP (as an indicator for the degree of out-of-distribution extrapolation required), we project matminer [53] features for all initial structures as well as relaxed MP structures onto two UMAP dimensions in fig. 10.
Finally, fig. 12(a) shows the elements-per-structure distribution of MP, MPtrj and WBM, normalized by dataset size. The mode of all three datasets is 3, but WBM’s share of ternary phases is noticeably more peaked than MP’s, which includes small numbers of unary and senary phases. The histogram in fig. 12(b) shows the distribution of the number of sites in MPtrj structures. The inset displays the same histogram log-scaled y-axis as well as a cumulative line to show that 90% of MPTrj structures contain fewer than 70 sites.
A.11 Predicted Formation Energy Analysis
We show the distribution for all materials in the WBM test set of model-predicted formation energies vs DFT ground truth in fig. 17 and the distribution of convex hull distance errors projected onto elements by composition for MACE and CHGNet in fig. 18.

### v3

@@ 4 Discussion
We have demonstrated the effectiveness of ML-based triage in high-throughput materials discovery and posit that the benefits of including ML in discovery workflows now clearly outweigh the costs. LABEL:{tab:metrics-table-uniq-protos} shows that in a realistic benchmark scenario that several models achieve a discovery acceleration greater than 2.5 across the whole dataset and up to 6 when considering only the 10k most stable predictions from each model (table S2). When starting this project, we were unsure which is the most promising ML methodology for high-throughput discovery. Our findings demonstrate a clear superiority in the accuracy and extrapolation performance of UIPs. Modeling forces enables these models to chart a path through atomic configuration space closer to the DFT-relaxed structure from where a more informed final energy prediction is possible.
As the convex hull becomes more comprehensively sampled through future discoveries, the fraction of unknown stable structures will naturally decline. This will lead to less enriched test sets and, consequently, more challenging and discriminative discovery benchmarks. However, the discovery task framed here addresses only a limited subset of potential UIP applications. We believe that additional benchmarks are essential to effectively guide UIP development. These efforts should prioritize task-based evaluation frameworks that address the four critical challenges we identify for narrowing the deployment gap: adopting prospective rather than retrospective benchmarking, tackling relevant targets, using informative metrics, and scalability.
Looking ahead, the consistently linear log-log learning curves observed in related literature [vonlilienfeld_retrospective_2020] suggest that further decreases in the error of UIPs can be readily unlocked with increased training data. This has be borne out in the scaling results of GNoME [merchant2023scaling], MatterSim [yang2024mattersim], Alexandria [schmidt2024improving], and OMat24 [barroso2024open] which all show improvements in performance when training on much larger datasets. To realize the full potential of scaling these models, future efforts should deploy their resources to generate large quantities of higher-than-PBE fidelity training data. The quality of a UIP model is circumscribed by the quality and level of theory of its training data.
@@
Beyond simply predicting thermodynamic stability at zero Kelvin, future models will need to understand and predict material properties under varying environmental conditions, such as finite temperature and pressure, to aid in materials discovery. In this context, temperature-dependent dynamical properties constitute an area ripe for interatomic potentials. Another key open question is how effectively these models can contribute to the computational prediction of synthesis pathways. Many existing approaches for reaction pathway prediction involve the use of heuristic rules for dealing with the significant added complexity of metastability alongside traditional ground state ab-initio data [mcdermott_graph-based_2021, aykol_rational_2021, wen_chemical_2023]. These algorithms will massively benefit from more efficient estimates of reaction energy barriers [yuan2024analytical] and non-crystalline, out-of-equilibrium materials [aykol_thermodynamic_2018], opening up a whole new field to machine learning accelerated inquiry.
@@ 5 Acknowledgments
We thank Rickard Armiento, Felix A. Faber and Abhijith S. Parackal for helping develop the evaluation procedures for Wren upon which this work builds. We also thank Rokas Elijosius for assisting in the initial implementation of Wrenformer and Mark Neumann, Luis Barroso-Luque and Yutack Park for submitting compliant models to the leaderboard.
@@ 6 Author Contributions
Janosh Riebesell: Methodology, Software, Data Curation, Investigation (Training: CGCNN, CGCCN+P, Wrenformer, Voronoi RF), Validation, Formal Analysis, Writing – original draft. Rhys Goodall: Conceptualization, Software, Validation, Formal Analysis, Writing – original draft, review & editing. Philipp Benner: Software, Investigation (Training: ALIGNN, MACE), Writing – original draft. Yuan Chiang: Investigation (Training: MACE), Formal Analysis, Writing – review & editing. Bowen Deng: Data Curation (MPtrj), Investigation (Training: CHGNet), Writing – review & editing. Gerbrand Ceder: Supervision, Funding Acquisition. Mark Asta: Supervision, Funding Acquisition. Alpha Lee: Supervision. Anubhav Jain: Supervision. Kristin Persson: Supervision, Writing – review & editing, Funding Acquisition.
@@ 8 Data availability
We chose the latest Materials Project (MP) [jain_commentary_2013] database release (v2022.10.28 at time of writing) as the training set and the WBM dataset [wang_predicting_2021] available at https://figshare.com/articles/dataset/22715158 as the test set for this benchmark. A snapshot of every ionic step including energies, forces, stresses and magnetic moments in the MP database is available at https://figshare.com/articles/dataset/23713842.
@@ Appendix A Supplementary Information
Note S1.1 Metrics on full test set and for 10k materials predicted most stable
@@
Unlike table 1 which evaluates model performance on the subset of unique WBM protostructures, table S1 includes the full WBM test set of 257k materials. The 44k additional materials in table S1 excluded in table 1 comprise discarded due to having a matching protostructure in MP plus another materials which are protostructure duplicates of another WBM material with lower energy. The most noteworthy difference between the two tables is a drop in DAF for all models in table S1 compared to table 1. MACE for example achieves a DAF of 3.5 on the full test set (table S1) compared to 3.85 on the subset of 215.5k materials with unique protostructures (table 1). This 10% increase is largely due to a 10% decrease in the fraction of materials below the MP convex hull: 15.3% (32,942 out of 215,488) in table 1 vs 16.7% (42,825 out of 256,963) in the full dataset (table S1). Since DAF is the ratio of the model’s precision for stability prediction to the prevalence of stable structures, a lower prevalence results in a higher DAF.
@@
While protostructures are non-trivial to match, thus potentially introducing bias into the test set by trying to deduplicate them, we still opted to feature table 1 in the main text since the removal of overlapping protostructures with MP makes it more closely reflect a model’s true ability to extrapolate to out-of-domain materials. Protostructure assignment is based on a combination of Aflow-style prototype labels and the chemical system [hicks_aflow_2021] we use the get_protostructure_label_from_spglib implemented in aviary [goodall2022aviary]. This string encodes the crystal’s prototype and chemical system and is invariant to symmetry preserving changes to the structure. Consequently structures with different lattice parameters are flagged as protostructure duplicates as they would both be expected to relax to the same ground state in experimental conditions.
@@
A real-world discovery campaign is unlikely to validate all stable predictions from a given model as we did in LABEL:{tab:metrics-table-uniq-protos}. Presumably, it will rank model predictions from most to least stable and follow that list as far as time and compute budget permits. Assuming that increases in compute resources will allow average future discovery campaigns to grow in scope, we believe 10k model validations to be a reasonable scope for average campaigns. This is what table S2 simulates by calculating classification and regression metrics for the 10k test set materials predicted to be most stable by each model. We again show dummy performance in the bottom row. Note that each model is now evaluated on a different slice of the data. However, the bottom row still shows dummy performance across the whole dataset.
Note S1.2 ROC Curves
@@
A material is classified as stable if the predicted is below the stability threshold. Since all models predict , they are insensitive to changes in the threshold . The receiver operating characteristic (ROC) curve for each model is plotted in fig. S1. The diagonal ‘No skill’ line shows the performance of a dummy model that randomly ranks material stability.
Note S1.3 Parity Plots
@@
Figure S2 shows that all models do well for materials far below the convex hull (left side of the plot). Performance for materials far above the convex hull is more varied with occasional underpredictions of the energy of materials far above the convex hull (right side). All models suffer most in the mode of the distribution at .
@@
Wrenformer is the only model with a large number of severe energy overpredictions at along the positive axis. We investigated these failure cases in more detail and found these overpredictions to be dominated by spacegroup 71 with poor representation in the training data. Digging into the MP-crystals-2022.10.28 dataset we see that on post analysis there are a large number of A2BC compounds in spacegroup 71 in the v2022.10.28 that are trapped in local minima and thus much higher in energy. Given that Wrenformer nominally predicts the energy the global minima conditioned on the protostructure these are unsuitable for training Wrenformer and their inclusion leads to the errors seen on similar, correctly relaxed structures, in the WBM test set.
@@
Beyond these MACE outliers visible in the plot, MACE exhibits another rare but reproducible type of failure case, in which the final predicted energy after relaxation is off by several orders of magnitude. The largest ‘derailed’ prediction was eV/atom for wbm-3-31970 (formula H2Ir). In each case, the MACE relaxation exhausted the maximum number of ionic steps set to 500 and caused a volume implosion from initial cell volumes of hundreds to relaxed cell volumes of tens of cubic Angstrom. Using the checkpoint trained on the M3GNet dataset which we received from the MACE authors, this failure mode occurred for several hundred of the 250k test set crystals. Using the checkpoint we trained ourselves on the MPtrj dataset, it affects only 44 test crystals, suggesting that these holes in the MACE PES can perhaps be fully plugged by further increasing the training set or even changing the loss function. Further analysis is ongoing. Since these derailed values are easily identified in practice when actually performing a prospective discovery campaign, we excluded them from the MACE parity plat and all other downstream analyses.
Note S1.4 Hull Distance Box plot
Figure S4 shows a box-plot of the errors in predicting the distance to the convex hull. The box demarks the interquartile range and the whiskers show the 5th and 95th percentiles. BOWSR has the largest median error, while Voronoi RF has the largest IQR. Looking at the models we see that there is a systematic trend to underpredict the distance to the convex hull. This trend can be rationalized give the bias in the Materials Project towards stable and therefore low formation energy structures. In contrast, due to it’s construction the WBM test set is distributionally distinct with a higher average formation energy. Therefore we propose that models trained on the Materials Project are likely result in such systematic underpredictions and this affect should be able to be addressed by training on additional data from a similar distribution to WBM i.e. Alexandria [schmidt2024improving].
Note S1.5 Classification Histograms using Model-Predicted Energies
Figure S5 shows histograms of the model predicted distances to the convex hull colored by the stability classification. These plots allow for practitioners to assess how the accuracy varies for materials depending on how far above or below the hull they are predicted to lie. The most pronounced affect we see from the histograms is that the CGCNN+P histogram is more strongly peaked than CGCNN’s giving much better agreement with the actual DFT ground truth distribution of hull distances for the test set. This explains why CGCNN+P performs better as a regressor, but also reveals how it can perform simultaneously worse as a classifier. By moving predictions closer to the stability threshold at above the hull, even small errors are significant enough to tip a classification over the threshold.
Note S1.6 Measuring extrapolation performance from WBM batch robustness
@@
Figure S6 shows the rolling MAE as a function of distance to the convex hull for each of the 5 WBM rounds of elemental substitution. These plots show a stronger performance decrease for Wrenformer and Voronoi RF than for UIPs and even force-less GNNs with larger errors like ALIGNN, MEGNet and CGCNN.
Figure S6 shows the rolling MAE for different models split out by calculation batch in the WBM data set. MEGNet and CGCNN both incur a higher rolling MAE than Wrenformer across all 5 batches and across most or all of the hull distance range visible in these plots. However, similar to the UIPs, MEGNet and CGCNN exhibit very little degradation for higher batch counts. The fact that higher errors for later batches are specific to Wrenformer and Voronoi RF suggests that training only on composition and coarse-grained structural features (spacegroup and Wyckoff positions in the case of Wrenformer; coordination numbers, local environment properties, etc. in the case of Voronoi RF) alone is insufficient to learn an extrapolatable map of the PES.
@@
Note S1.7 Largest Errors vs. DFT Hull Distance
@@
Given the large variety of models tested, we asked whether any additional insight into the errors can be gained from looking at how the predictions vary between different models. In fig. S7 we see two distinct groupings emerge when looking at the 200 structures with the largest errors. This clustering is particularly apparent when points are colored by model disagreement.
Note S1.8 Exploratory Data Analysis
To give high-level insights into the MP training and WBM test set used in this work, we include element distributions for structures in both datasets (fig. S8). To show how frame selection from MP structure relaxation affected relative elemental abundance between MP relaxed structures and the snapshots in MPtrj, fig. S9 shows element occurrence ratios between MPtrj and MP. Similarly, fig. S10 shows the elements-per-structure distribution of MP, MPtrj and WBM, normalized by dataset size. The mode of all three datasets is 3, but WBM’s share of ternary phases is noticeably more peaked than MP’s, which includes small numbers of unary and senary phases. fig. S11 plots the distributions of formation energies, decomposition energies and convex hull distances (with respect to the convex hull spanned by MP materials only) for the MP and WBM data sets. This plot not only gives insight into the nature of the dataset but also emphasizes the increased difficulty of stability vs formation energy prediction arising from the much narrower distribution of convex hull distances compared to the more spread-out formation energy distribution. For WBM we see that the formation energy distribution exhibits much wider spread than the convex hull distance distribution, spanning almost vs less than spread in the hull distances. This highlights why stability prediction is a much more challenging task than predicting energy of formation. It requires correctly ranking the subtle energy differences between chemically similar compounds in the same chemical system rather than comparing a single material with the reference energies of its constituent elements. DFT has been shown to significantly benefit from the systematic cancellation of errors between chemically similar systems when trying to identify the lowest-lying polymorph [hautier_accuracy_2012]. This beneficial cancellation has yet to be conclusively demonstrated for ML stability predictions. So far, only the lack thereof has been shown in [bartel_critical_2020] where they encountered a much more random error distribution among similar chemistries than simulations from first principles.
Finally, Believing MPtrj to be an influential dataset for the near-term continued development of universal interatomic potentials, we plot histograms showing the distributions of target values for energies, forces, stresses and magnetic moments in fig. S12. The histogram in fig. 12(e) shows the distribution of the number of sites in MPtrj structures. The inset displays the same histogram log-scaled y-axis as well as a cumulative line to show that 90% of MPtrj structures contain fewer than 70 sites. fig. S13 shows the forces and magnetic moments for the MPTrj dataset projected onto different elements.